# Catastrophic Liability: Managing Systemic Risks in Frontier AI Development

### v1

@@ 3.3 Healthcare
Negative consequences of AI use in healthcare do not clearly fit within this liability framework. The Stanford University Human-Centered Artificial Intelligence Institute’s brief (Mello & Guha, 2024) on liability for AI’s use in the healthcare sector underscores as a key problem the inability of the plaintiff to identify breaches of “duty of care” on the part of healthcare providers versus malfunctions of specific components of the AI system. Additionally, it is difficult for the healthcare providers to identify appropriate uses of the systems. The recommendation within the brief is to have “disclosure requirements” given to physicians and hospitals. Such disclosures particularly focus on model performance, particularly for different demographics of patients, and are intended to enable transparency in the decision making of the healthcare providers. Although we believe the disclosure requirements are a step in the right direction, we also believe that more transparency in terms of development is needed in order to ensure safe use of AI in Healthcare. The transparency in both development and use of the system should allow Product Liability protection.
The healthcare sector’s experience with AI liability offers several crucial lessons for broader AI development policy:
First, it demonstrates how opacity in AI systems can create fundamental challenges for establishing liability. When adverse outcomes occur, the inability to clearly attribute responsibility between human operators, AI systems, and development practices creates significant legal and practical obstacles. This suggests that, absent strict liability, frontier AI development requires comprehensive documentation of both system capabilities and testing procedures to enable meaningful accountability.
Second, the focus in healthcare on cross-demographic performance differences highlights the importance of thorough testing across diverse scenarios. For frontier AI systems, this principle suggests the need for extensive documentation of model behavior across different contexts and potential use cases, not just aggregate performance metrics.
Third, the healthcare sector’s emphasis on disclosure requirements for physicians offers a model for how transparency obligations could be structured for AI developers. Just as healthcare providers need clear information about AI system limitations to make informed decisions, stakeholders interacting with frontier AI systems need detailed understanding of system capabilities and constraints. These lessons suggest AI development policy should mandate:
@@
The healthcare sector shows that effective liability frameworks require transparency in both development and deployment phases. Without visibility into development practices, even well-documented deployment problems may not lead to appropriate accountability.
@@ 4.1 EU AI Act and AI Liability Directive
The EU AI Liability Directive, which would have complemented the EU AI Act and the Product Liability Directive, was proposed in 2022 and scrapped in 2025. The EU AI Act establishes documentation and reporting requirements for both high-risk AI systems and general-purpose AI models with systemic risk (European Parliament, 2024).
For high-risk systems, the AI Liability Directive would have granted plaintiffs access rights to documentation of system functioning and usage to establish duty of care (Article 3). Required documentation included accuracy metrics, robustness measures, cybersecurity protocols, input data specifications, and system performance data across user groups (Article 13-14) (European Commission, 2022). In the absence of this directive, it is unclear whether plaintiffs will have access to this information.
@@
Article 3 establishes that national courts can compel AI providers to disclose evidence about suspected damages from high-risk AI systems. If providers fail to comply with disclosure orders, courts will presume non-compliance with duty of care requirements, though providers retain the right to rebut this presumption (European Parliament, 2024).
Since today’s most advanced AI models (as of early 2025) can already be foreseeably misused to disseminate illegal, false, or discriminatory content, or to disrupt democratic processes (Kreps & Kriner, 2023), the EU AI Act appears to treat current frontier AI systems as posing a non-negligible level of systemic risk. However, the operational methodology for determining “high-impact capabilities” has not yet been defined. This is important to resolve before the Article 111(3)’s compliance deadline of August 2nd, 2027, for providers of general-purpose AI models placed on the market before August 2nd, 2025. If a reasonable interpretation of the AI Act implies that today’s models pose systemic risks, then frontier labs should act urgently to comply with Section 3’s “Obligations of providers of general-purpose AI models with systemic risk.” Providers of general-purpose AI models with systemic risk face additional documentation requirements beyond standard model cards, including:
@@
    System architecture explanations detailing component integration (European Parliament, 2024, Annex XI, Section 2).
@@
    Implementing alternative adequate measures if standards are unavailable (European Parliament, 2024, Recital 117).
The AI Office will facilitate the development of codes of practice in collaboration with civil society organizations and experts to establish risk taxonomies and mitigation measures (European Parliament, 2024, Recital 116).
@@
As summarized in Section 3.1, the duty of care standard in U.S. tort law heavily relies on industry standards and established practices. The artificial intelligence risk management framework (AI RMF) released by the U.S. National Institute of Standards and Technology can be used as a basis for these standards. The AI RMF is entirely voluntary, and NIST emphasizes that practitioners are free to apply the framework’s core functions “as best suits their needs for managing AI risks based on their resources and capabilities” (Tabassi, 2023). This begs further clarification, some of which is available in the Generative Artificial Intelligence Profile from 2024 (National Institute of Standards and Technology (2024), US), but neither document gives a clear template for governance and reporting requirements in frontier AI development. That said, since frontier AI labs are vocal about their perceived risks from AI, and since the compute requirements for AI with systemic risks carry high costs, it is safe to assume that frontier AI development furnishes the needs, resources, and capabilities for most or all of the framework to apply.
Other resources that could be used to determine the standards and practices of AI development in the U.S. include the Voluntary AI Commitments from AI companies secured by the Whitehouse in 2023 (House, 2023) and the Partnership on AI (PAI) report Risk Mitigation Strategies for the Open Foundation Model Value Chain (Srikumar et al., 2024). The actual practices in AI development, whether learned solely by collecting publicly available information, revealed by audits, or obtained via subpoena, would likely also be considered in a tort case.
@@
Unfortunately, many frontier AI labs do not appear to follow GV-1.3-007 (Stein-Perlman, 2024), creating a gap between industry “practices” and “standards” that may still constitute a breach of their duty of care. Given the precedent from the 1932 tugboat case, the level of care considered appropriate by courts may prioritize “reasonable” measures over common practice. Presumably, practices described in the AI RMF would be considered reasonable, especially given NIST’s even-handed reputation and the stakeholder input process used in the creation of the RMF. Thus, unless their policies on the matter exist but are not available to the public, it appears that many AI companies are already breaching at least one duty of care.
We encourage such companies to take immediate actions to adhere more closely to the RMF’s recommendations, in order to prevent potential harms from excessively risky AI and to reduce their own liability for any harms that occur.
@@ 5 Discussion
Our examination of liability frameworks reveals both the promise and challenges of applying existing safety documentation practices to Frontier AI development. While the industries we reference offer valuable precedents for managing catastrophic risks through organizational governance and independent oversight, the novel features of general-purpose AI systems necessitate expanding traditional concepts of duty of care and liability.
5.1 Key Recommendations
Based on our analysis of nuclear energy, aviation, and healthcare liability frameworks; and current U.S. and EU regulatory developments, we propose the following recommendations organized by stakeholder groups.
5.1.1 Policymakers and Regulators
  1. 1.
    Establish clear liability frameworks that:
    1. (a)
      Define duty of care based on reasonable safety practices rather than just industry customs;
    2. (b)
      Create mechanisms for attributing responsibility across complex AI development pipelines;
    3. (c)
      Address both high-risk applications and systemic risks from general-purpose AI.
  2. 2.
    Implement certification requirements similar to the Price-Anderson Act, including:
    1. (a)
      Federal certification of safety practices;
    2. (b)
      Insurance-based liability caps for certified developers;
    3. (c)
      Clear chain of responsibility for downstream harms.
  3. 3.
    Develop specific guidance for complying with EU AI Act requirements, particularly:
    1. (a)
      Documentation requirements for general-purpose AI models with systemic risk;
    2. (b)
      Protocols for demonstrating compliance with proportional obligations;
    3. (c)
      Standards for evaluating high-impact capabilities.
5.1.2 Standards Organizations
  1. 1.
    Strengthen the NIST AI Risk Management Framework by:
    1. (a)
      Requiring evidence of active engagement with safety challenges;
    2. (b)
      Mandating processes to prevent “strategic overlooking” of risks;
    3. (c)
      Providing detailed templates for governance practices;
    4. (d)
      Requiring transparent documentation of risk trade-off decisions;
  2. 2.
    Define robust third-party audit mechanisms that would:
    1. (a)
      Have authority to access and evaluate safety documentation;
    2. (b)
      Maintain confidentiality of trade secrets while ensuring accountability;
    3. (c)
      Assess novel risks that emerge during development;
    4. (d)
      Include diverse expertise and stakeholder perspectives.
5.1.3 Frontier AI Developers
  1. 1.
    Implement comprehensive documentation of safety measures throughout the AI development process, including:
    1. (a)
      Risk assessment methodologies and results;
    2. (b)
      Testing protocols and outcomes across diverse scenarios;
    3. (c)
      Decision-making processes for safety-critical choices;
    4. (d)
      Organizational responses to identified risks;
    5. (e)
      Plans for halting development if unacceptable risks emerge.
  2. 2.
    Establish clear standards for demonstrating meaningful human oversight, including:
    1. (a)
      Explicit delineation of human versus AI system responsibilities;
    2. (b)
      Documentation of oversight mechanisms and their effectiveness;
    3. (c)
      Regular assessment of human operator capabilities relative to system complexity.

### v2

@@ 3.3 Healthcare
Negative consequences of AI use in healthcare do not clearly fit within this liability framework. The Stanford University Human-Centered Artificial Intelligence Institute’s brief (Mello and Guha 2024) on liability for AI’s use in the healthcare sector underscores as a key problem the inability of the plaintiff to identify breaches of ”duty of care” on the part of healthcare providers versus malfunctions of specific components of the AI system.
Additionally, it is difficult for the healthcare providers to identify appropriate uses of the systems. The recommendation within the brief is to have ”disclosure requirements” given to physicians and hospitals. Such disclosures particularly focus on model performance, particularly for different demographics of patients, and are intended to enable transparency in the decision making of the healthcare providers. Although we believe the disclosure requirements are a step in the right direction, we also believe that more transparency in terms of development is needed in order to ensure safe use of AI in Healthcare. The transparency in both development and use of the system should allow Product Liability protection.
Implications.
The healthcare sector’s experience with AI liability offers a critical lesson for frontier AI development: even in high-risk applications with clear human oversight, opacity in AI systems fundamentally undermines accountability. This makes the case for transparency even more compelling when considering frontier AI’s systemic risks:
First, healthcare demonstrates that attribution challenges exist even when AI systems operate under direct human supervision with clear intended uses. When adverse outcomes occur in a medical setting—where physicians actively review AI recommendations and maintain ultimate decision-making authority—courts still cannot determine whether harm resulted from physician error or AI malfunction. This inability to establish causation in supervised, domain-specific applications suggests that frontier AI systems, which may operate autonomously across multiple domains, will present exponentially greater attribution challenges.
Second, the focus in healthcare on cross-demographic performance differences highlights a critical gap: current AI systems can fail in ways that are invisible until deployment across diverse populations. While healthcare AI affects specific patient groups, frontier AI systems could exhibit similar hidden failure modes that affect entire societies. The healthcare sector’s struggle to identify these failures even with extensive clinical oversight underscores why comprehensive development documentation becomes essential for systems with systemic risk potential.
Third, the healthcare sector’s emphasis on disclosure requirements reveals an important limitation: transparency at deployment is insufficient when the root causes of failure lie in development choices. Healthcare providers cannot assess ”appropriate uses of the systems” without understanding how they were built and tested. For frontier AI, where inappropriate use could have catastrophic consequences, this gap between deployment transparency and development opacity becomes potentially existential.
These lessons suggest AI development policy should mandate:
@@
The healthcare sector shows that effective liability frameworks require transparency in both development and deployment phases. Without visibility into development practices, even well-documented deployment problems may not lead to appropriate accountability. If attribution is this challenging in healthcare’s controlled environment, the urgency for comprehensive documentation in frontier AI development—where systems may operate beyond human oversight—becomes undeniable.
3.4 Cybersecurity and Open Source Software
The cybersecurity domain offers crucial lessons for AI liability frameworks, particularly in distinguishing between high-risk applications and systemic risks from general-purpose models. While traditional security practices work well for specific AI deployments, frontier AI development requires fundamentally different approaches to transparency and accountability.
Security challenges across risk levels.
For AI in high-risk applications (like healthcare diagnostics or hiring systems), cybersecurity’s Common Vulnerabilities and Exposures (CVE) model provides clear benefits: standardized vulnerability reporting, coordinated patching, and defined responsibilities (Cattell, Ghosh, and Kaffee 2024). However, general-purpose AI models with systemic risk present novel challenges. Unlike traditional software where vulnerabilities can be patched, AI models can develop emergent capabilities during training that persist throughout deployment. A text generation model might unexpectedly acquire the ability to assist with cyberattacks or bioweapon design—risks that cannot be simply ”patched” without full retraining.
This distinction parallels the nuclear industry’s approach discussed in Section 4.1: while specific safety violations can be addressed through targeted fixes, systemic risks require comprehensive documentation throughout development. For frontier AI, this means documenting not just final capabilities but the entire training process, including unexpected behaviors discovered during development.
Distributed responsibility for systemic risks.
The cybersecurity principle of distributed responsibility takes different forms across risk levels. For high-risk applications, the CVE system’s clear attribution works well—a vulnerability in a medical AI system can be traced to specific code or training data. But for general-purpose models with systemic risk, responsibility becomes more complex. When a foundation model develops dangerous capabilities, liability may span data providers who supplied training materials, compute providers who enabled large-scale training, and developers who chose training objectives.
Industry responses like secure model formats (Abid Ali Awan 2023; Boykis 2024) demonstrate how technical standards can help, but only partially. While these formats prevent code injection attacks, they cannot address the deeper challenge of models that have learned harmful capabilities. This suggests liability frameworks must distinguish between security measures appropriate for deployment (mainly high-risk and domain-dependent) versus development (mainly systemic risk).
Open source and innovation.
Despite recent political reticence towards ”safety” framings (Sivaram 2025), innovation-minded stakeholders need not shy away from risk management; in fact, reducing liability can advance innovation rather than hinder it. Clear liability frameworks provide certainty that enables responsible developers to invest confidently in frontier AI research, knowing they won’t face unpredictable legal exposure for following established safety practices (Merwe, Ramakrishnan, and Anderljung 2024). This parallels how the Price-Anderson Act’s liability caps actually encouraged nuclear power development by providing predictable risk allocation.
The tension between open science collaboration and dual-use AI risks requires nuanced approaches that preserve innovation incentives while managing systemic risks. ”Openness” in AI helps democratize access to AI capabilities and enables crucial safety research through transparency (White et al. 2024). However, it also presents challenges when models develop unexpected dangerous capabilities that cannot be easily recalled or patched (Al-Kharusi et al. 2024). Responses to open source AI must distinguish between different types of open source contributions—rewarding transparency and collaborative safety research while establishing clear responsibilities for developers who release models with predictably risky capabilities (Sidhpurwala et al. 2024). We need to balance the needs for open science and collaboration with the risks of widespread access to dual-use AI (NTIA 2024) and autonomous AI that poses systemic risks even when deployed for benign purposes (Mitchell et al. 2025).
Implications for liability frameworks.
These distinctions suggest different transparency requirements based on risk level:
For high-risk AI applications: The cybersecurity model largely suffices—standardized reporting, coordinated disclosure (Cattell, Ghosh, and Kaffee 2024), and clear attribution of vulnerabilities to specific components. These practices are sufficient because the scale and vectors of risk are alike.
For general-purpose AI with systemic risk, risks can emerge spontaneously during development (Taylor and Bergen 2025) or post-deployment (Yu et al. 2024; Hubinger et al. 2024). If misaligned, autonomous AI models are created, they may be difficult or impossible to shut down once released, leading to existential risks (Pan et al. 2025). Furthermore, algorithmic assessments may be unable to set upper bounds for AI capabilities and risks (Barnett and Thiergart 2024). Thus, transparency must extend throughout the development life-cycle:
  • Before-training documentation of intended capabilities and risk assessments (BigCode et al. 2023)
  • White-box audits of AI systems during development (Casper et al. 2024)
  • ”Pause and assess” protocols when unexpected capabilities emerge—similar to the NIST framework’s GV-1.3-007 discussed in Section 5.2
  • Collaborative reporting on emergent behaviors in deployed models (Cattell, Ghosh, and Kaffee 2024)
The cybersecurity domain thus demonstrates that while transparency and documentation benefit all stakeholders, the nature of required transparency fundamentally differs between high-risk applications and systemic risks. For the latter, traditional ”patch and fix” approaches are insufficient. Instead, liability frameworks must incentivize comprehensive documentation of the development process itself, creating accountability for risks that emerge before deployment—a crucial distinction for frontier AI governance.
@@ 4.1 EU AI Act and AI Liability Directive
The EU AI Liability Directive, which would have complemented the EU AI Act and the Product Liability Directive, was proposed in 2022 and scrapped in 2025. The EU AI Act establishes documentation and reporting requirements for both high-risk AI systems and general-purpose AI models with systemic risk (European Parliament 2024).
For high-risk systems, the AI Liability Directive would have granted plaintiffs access rights to documentation of system functioning and usage to establish duty of care (Article 3). Required documentation included accuracy metrics, robustness measures, cybersecurity protocols, input data specifications, and system performance data across user groups (Article 13-14) (European Commission 2022). In the absence of this directive, it is unclear whether plaintiffs will have access to this information.
@@
Article 3 establishes that national courts can compel AI providers to disclose evidence about suspected damages from high-risk AI systems. If providers fail to comply with disclosure orders, courts will presume non-compliance with duty of care requirements, though providers retain the right to rebut this presumption (European Parliament 2024).
Since today’s most advanced AI models (as of early 2025) can already be foreseeably misused to disseminate illegal, false, or discriminatory content, or to disrupt democratic processes (Kreps and Kriner 2023), the EU AI Act appears to treat current frontier AI systems as posing a non-negligible level of systemic risk. However, the operational methodology for determining “high-impact capabilities” has not yet been defined. This is important to resolve before the Article 111(3)’s compliance deadline of August 2nd, 2027, for providers of general-purpose AI models placed on the market before August 2nd, 2025. If a reasonable interpretation of the AI Act implies that today’s models pose systemic risks, then frontier labs should act urgently to comply with Section 3’s “Obligations of providers of general-purpose AI models with systemic risk.” Providers of general-purpose AI models with systemic risk face additional documentation requirements beyond standard model cards, including:
@@
    System architecture explanations detailing component integration (European Parliament 2024, Annex XI, Section 2).
@@
    Implementing alternative adequate measures if standards are unavailable (European Parliament 2024, Recital 117).
The AI Office will facilitate the development of codes of practice in collaboration with civil society organizations and experts to establish risk taxonomies and mitigation measures (European Parliament 2024, Recital 116).
@@
As summarized in Section 3.1, the duty of care standard in U.S. tort law heavily relies on industry standards and established practices. The artificial intelligence risk management framework (AI RMF) released by the U.S. National Institute of Standards and Technology can be used as a basis for these standards. The AI RMF is entirely voluntary, and NIST emphasizes that practitioners are free to apply the framework’s core functions “as best suits their needs for managing AI risks based on their resources and capabilities” (Tabassi 2023). This begs further clarification, some of which is available in the Generative Artificial Intelligence Profile from 2024 (National Institute of Standards and Technology (2024) US), but neither document gives a clear template for governance and reporting procedures in frontier AI development. That said, since frontier AI labs are vocal about their perceived risks from AI, and since the compute requirements for AI with systemic risks carry high costs, it is safe to assume that frontier AI development furnishes the needs, resources, and capabilities for most or all of the framework to apply.
Other resources that could be used to determine the standards and practices of AI development in the U.S. include the Voluntary AI Commitments from AI companies secured by the Whitehouse in 2023 (House 2023) and the Partnership on AI (PAI) report Risk Mitigation Strategies for the Open Foundation Model Value Chain (Srikumar, Chang, and Chmielinski 2024). The actual practices in AI development, whether learned solely by collecting publicly available information, revealed by audits, or obtained via subpoena, would likely also be considered in a tort case.
@@
Unfortunately, many frontier AI labs do not appear to follow GV-1.3-007 (Stein-Perlman 2024), creating a gap between industry “practices” and “standards” that may still constitute a breach of their duty of care. Given the precedent from the 1932 tugboat case, the level of care considered appropriate by courts may prioritize “reasonable” measures over common practice. Presumably, practices described in the AI RMF would be considered reasonable, especially given NIST’s even-handed reputation and the stakeholder input process used in the creation of the RMF. Thus, unless their policies on the matter exist but are not available to the public, it appears that many AI companies are already breaching at least one duty of care.
One aspect that might make the project of implementing such a liability framework seem doomed for failure in the U.S. is the current Trump administration’s focus on the US removing “barriers” to AI development rather than ensuring AI safety. This is particularly demonstrated in the House passing the “The Big Beautiful Bill” calling for a banning of state level regulation of AI development (119th Congress 2025). While it remains to be seen whether such legislation will go through, there are ways that such a framework might still be helpful.
For instance, the liability framework might still be beneficial even if it is ultimately voluntary. Additionally, there have been noted legal issues that can come with the administration attempting to enforce states from regulating (Novelli, Gaur, and Floridi 2025) as well as potential conflict with the administrations interest on “human flourishing.” For these reasons, it seems that developing an auditing framework is beneficial. Specifically, we encourage such companies to take immediate actions to adhere more closely to the RMF’s recommendations, in order to prevent potential harms from excessively risky AI and to reduce their own liability for any harms that occur.
@@ 5 Discussion
Our examination of liability frameworks from nuclear energy, aviation, healthcare, and cybersecurity demonstrates that comprehensive safety documentation creates competitive advantages while improving oversight—challenging the false dichotomy between innovation and accountability. The evidence supports three key conclusions that reshape how we should approach frontier AI governance.
Voluntary standards already carry legal force.
The assumption that voluntary AI commitments lack enforcement mechanisms fundamentally misunderstands U.S. tort law. As demonstrated in the 1932 tugboat case, courts determine negligence based on “reasonable practice” rather than common practice. This means courts may consider frameworks like NIST’s AI Risk Management Framework as evidence of reasonable practice, even without mandatory regulation. When frontier AI companies publicly acknowledge catastrophic risks and commit to safety standards, they strengthen the case that such standards represent reasonable care.
This legal reality means that many frontier AI labs may be exposed to negligence claims. The NIST AI RMF explicitly calls for “a plan to halt development or deployment of a GAI system that poses unacceptable negative risk” (GV-1.3-007), yet public evidence suggests many companies lack such protocols. Given that these companies have both acknowledged AI risks and committed to following NIST standards, their failure to implement basic safety measures could lead U.S. courts to find negligence under current law.
Systemic risks require expanded duties of care.
The EU AI Act’s distinction between high-risk applications and general-purpose models with systemic risk reflects a crucial insight: unlike traditional AI applications that create dangers through specific deployments, frontier models’ capabilities can pose threats regardless of intended use. This distinction, supported by our case studies, requires fundamentally different approaches to liability and oversight.
In particular, the healthcare sector’s attribution chal- lenges—even under direct human supervision—demonstrate why opacity becomes exponentially more problematic for AI with systemic risk potential. When courts cannot deter- mine causation even in controlled medical settings, fron- tier AI systems operating autonomously across multiple do- mains will present insurmountable accountability gaps with- out comprehensive development documentation.
Transparency benefits all stakeholders.
Our case studies consistently demonstrate that transparency reduces liability exposure for responsible developers while enabling meaningful oversight. Nuclear energy’s governance frameworks prevent “strategic overlooking” of risks while reducing liability uncertainty. Aviation’s rigorous documentation, though initially costly, provides regulatory certainty that enables safe innovation. Even cybersecurity’s coordinated disclosure benefits developers by providing time to patch vulnerabilities before public disclosure, reducing exploitation risk while maintaining user trust.
Current opacity in AI development creates long-term risks for all parties: companies face unpredictable legal exposure, regulators cannot verify safety claims, and society bears undocumented risks. While some firms may gain short-term competitive advantages from secrecy, the solution is not less accountability but better documentation that serves everyone’s long-term interests.
5.1 Concrete Recommendations
Based on our analysis, we propose mutually beneficial actions for each stakeholder group that align individual incentives with collective safety:
Stakeholder Priority Action Rationale Key Deliverable
Standards Organizations Strengthen NIST AI RMF with nuclear-inspired governance templates Creates legal safe harbor for compliant developers while establishing clear benchmarks for negligence Detailed templates for organizational learning, risk trade-off documentation, and pause-and-assess protocols
Establish independent audit mechanisms following aviation DO-178C model External validation reduces liability uncertainty and ensures genuine safety culture Certification process with legal weight that reduces insurance costs
Policymakers & Regulators Implement liability caps tied to safety certification (Price-Anderson model) Provides predictable liability for responsible developers while maintaining accountability for negligent actors Federal certification program that caps liability for compliant companies
Clarify EU AI Act’s systemic risk evaluation criteria Regulatory certainty enables compliance planning and consistent global standards Specific methodologies for determining “high-impact capabilities” before August 2027
Frontier AI Developers Document all capability discoveries and safety measures throughout development Liability protection through demonstrable due care plus internal safety improvements Safety cases for each major model release, including unexpected capabilities discovered during training
Create independent safety boards with real decision authority Prevents “strategic overlooking” patterns that preceded Fukushima while building public trust Governance structures that can halt development when risks exceed acceptable thresholds
Table 1: Priority Recommendations for AI Safety Documentation by Stakeholder
These recommendations recognize that effective AI governance must work with market incentives rather than against them. Companies that invest in comprehensive safety documentation can achieve reduced legal risk and greater regulatory certainty—potentially creating competitive advantages for responsible development, as seen in other safety-critical industries.
5.2 Implementation Strategy
The path forward requires coordinated action across stakeholders, building on existing momentum rather than waiting for comprehensive regulation:
Near-term Actions (Within 12 months): Frontier AI companies should implement GV-1.3-007 protocols immediately to reduce current legal exposure. NIST should release detailed governance templates incorporating lessons from nuclear safety. EU regulators should clarify systemic risk evaluation criteria to enable compliance planning.
Medium-term Goals (1-2 years): Establish pilot certification programs linking safety documentation to liability protection. Develop standardized audit procedures for frontier AI development. Create international coordination mechanisms for consistent standards.
Long-term Vision (2+ years): Full integration of safety documentation requirements with competitive advantages. Mature governance frameworks that scale with advancing AI capabilities. Global standards that enable responsible innovation while preventing race-to-the-bottom dynamics.
5.3 Addressing Potential Objections
“This will slow AI progress”: Our case studies provide nuanced evidence. While aviation’s documentation requirements may increase development time and costs, they ultimately enable sustainable innovation by providing regulatory certainty and public trust. Likewise, safety requirements in nuclear energy have constrained some innovation while enabling the industry’s continued operation. For AI, comprehensive safety documentation may slow initial development but creates conditions for sustainable, trusted deployment.
“Voluntary standards are unenforceable”: U.S. tort law considers voluntary standards as evidence of reasonable practice, particularly when companies publicly commit to them. While not automatically binding, courts can and do use such standards to evaluate negligence claims. Companies cannot acknowledge catastrophic risks, commit to safety standards, and then ignore those commitments without potential legal consequences.
“Technical opacity makes documentation impossible”: Our cybersecurity analysis shows that different types of transparency serve different risk profiles. While complete interpretability remains challenging, documenting development processes, capability discoveries, and safety decisions is both feasible and legally protective.